OpenStudy (anonymous):

Factor by grouping xy - 3x - 8y + 24

OpenStudy (turingtest):

(x-8)(y-3)

OpenStudy (anonymous):

how did u get that turing test?

OpenStudy (anonymous):

xy - 3x - 8y + 24 = y(x-8) -3(x-8) \[=(x-8)(y-3)\]

OpenStudy (anonymous):

hope that helps.

OpenStudy (anonymous):

thank u i actually have 2 more cud u help me with them?

OpenStudy (anonymous):

4x^2 + 20x + 5y +xy

OpenStudy (turingtest):

hmmm not quite the same I'll have to think about it....

OpenStudy (anonymous):

its ok i got it

OpenStudy (turingtest):

(4x+y)(x+5)

OpenStudy (anonymous):

4x^2 + 20x + 5y +xy = 4x(x+5) +y(x+5) =(x+5)(4x+y)

OpenStudy (turingtest):

i just make the parenthesis first, then put in the signs. after that you can see the which numbers go where by looking at the coefficients

OpenStudy (turingtest):

probably not the way teachers show it though...